Activity: Amortization Payment

Cedric purchased a new fishing boat for

P130,000. He made a P20,000 down payment, and financed the balance at his bank

for 7 years. What amortization payments are required every 3 months, at 16%

interest, to pay off the boat loan?

2.Cameron Manufacturing recently purchased a new

computer system for P150,000. What amortization payment is required each month,

at 12% interest, to pay off this obligation in 8 years?

3The Clintons bought a home for P12,050,000.

After a 15% down payment, the balance is financed at 8% interest for 9 years.

(a) What equal quarterly payments will be required to amortize this mortgage

loan? (b) What is the total amount of interest the Clintons will pay on the

loan?


1
Expert's answer
2021-05-24T03:46:03-0400

1.

Loan amount(PV)=Price of a new fishing boat - Down payment

"=130,000-20,000\\=P110,000"


"Rate=\frac{Interest\space rate}{ Periods \space of\space compounding}"


"=\frac{16\%}{4}"


"=4\%"


Number of periods (Nper) = Loan term "\times" Periods of compounding

"=7\times 4\\=28"

We can compute the periodic payments by using the PMT function in Excel. The PMT function can be used as follows:

"=PMT(rate,nper,pv,[fv],[type])\\

=PMT(4\%,28,-110000)\\

=P6,601.43"


2.

Present value (PV)=P150,000

"Rate=\frac{Interest\space rate}{ Periods \space of\space compounding}"


"=\frac{12\%}{12}"


"=1\%"


Number of periods (Nper) = Loan term "\times" Periods of compounding

"=8\times12\\=96"

We can compute the periodic payments by using the PMT function in Excel. The PMT function can be used as follows:

"=PMT(rate,nper,pv,[fv],[type])\\

=PMT(1\%,96,-150000)\\

=P2,437.93"


3.

(i)

Loan amount(PV)=Price of a home - Down payment

"=12,050,000-(15\%\times12,050,000)\\=P10,242,500"

"Rate=\frac{Interest\space rate}{ Periods \space of\space compounding}"


"=\frac{8\%}{4}"


"=2\%"


Number of periods (Nper) = Loan term "\times" Periods of compounding


"=9\times 4\\=36"


We can compute the periodic payments by using the PMT function in Excel. The PMT function can be used as follows:

"=PMT(rate,nper,pv,[fv],[type])\\

=PMT(2\%,36,-10,242,500)\\

=P401,842.49"

(ii)

Total amount of payments = Quarterly payments "\times" Number of periods

"=401,842.49\times 96\\=P14,466,329.74"


Total amount of interest = Total amount of payments - loan amount

"=14,466,329.74-10,242,500\\=P4,223,829.74"